# Uncovering Reliable Indicators: Improving IoC Extraction from Threat Reports

### v1

@@ 3.3 Custom User Interface
The user interface provides the analyst with the LANCE-generated label and justification. The analyst has the option to either accept or override the generated label. The interface also enables analysts to view justifications by hovering over indicators as well as adding additional comments for future reference. After all labels are finalized by the analyst, they can be exported in a structured format along with all justifications and additional comments optionally provided by the analyst.
@@ 4 PRISM Dataset Creation
We now describe how we use the HITL system described earlier (Figure 1) to create our PRISM dataset. The labeling pipeline we use is depicted in Figure 3. We start by describing our ground truth generation methodology with human vetting, and then discuss our concrete application of this methodology to the PRISM dataset. We describe PRISM first here, and later (in Section 5) use the manually verified IoC labels obtained during this dataset creation process to evaluate LANCE’s accuracy by measuring its level of agreement with expert threat analysts.
@@ 4.1 Methodology Overview
To further increase the quality of the PRISM dataset as well as to evaluate the effects LANCE had on the junior analysts, we also employed a senior analyst. The senior analyst annotated only the reports that had indicators that the junior analysts disagreed on. This way, we minimized the time the senior analyst needed to spend on labeling indicators and used their expertise only on the difficult-to-label indicators.
@@ 4.1.1 Baseline Annotation Pass (BAP)
The reports that included at least one disputed indicator were tagged as disputed reports and were given to a senior analyst to determine the correct label of the indicators. These reports were given to the senior analyst through the same UI and with the same instructions as it was given to the junior analysts.
The senior analyst was tasked to label all indicators in the disputed reports, not only the disputed indicators. This way, we could see if there were other indicators that were labeled falsely by both the junior analysts and the LLM.
@@ 4.1.2 Guided Annotation Pass (GAP)
For the second part of the ground truth creation setup, we used the second set of reports. The setup for GAP was similar to that of BAP with the exception that the analysts did have access to the labels that were assigned to the indicators by LANCE, as well as the justifications that LANCE provided for this choice. This allowed the junior analysts to be “advised” by LANCE while labeling each indicator.
@@ 4.2 PRISM Dataset Details
To ensure relevance and quality, we filtered for reports that also appeared in AlienVault [7], a widely used threat intelligence platform. Specifically, we selected reports published by AlienVault’s official user account, which aggregates high-quality reports from vendors such as Kaspersky, Palo Alto Networks, and Microsoft. From this filtered corpus, we identified 50 reports that (i) appeared on both ORKL and AlienVault, (ii) were fewer than 30 pages in length, and (iii) contained more than 5 unique indicators. Although this threshold provides only an upper bound on the number of true IoCs, it serves as an effective proxy for report richness.
@@
TABLE II: Analysis of non-unique indicator instances in reports for BAP and GAP, including total sums and per IoC type.
Disagreements between annotators, or between annotators and the LANCE-generated labels, were flagged as disputed. We identified 15 reports from BAP that had at least one disputed indicator and 9 from GAP. These reports were escalated to the senior analyst, who re-labeled all indicators in those reports to ensure consistency and completeness. Figure 4 shows the agreement ratios between annotators and with LANCE.
@@
Figure 5: Comparison of LANCE and analyst performance across precision, recall, and F1 score during both annotation phases. The top row shows results from BAP (manual labeling without LANCE assistance), while the bottom row shows GAP (LANCE-assisted annotation).
@@ 5 IoC Labeling Evaluation
In this section, we evaluate LANCE by comparing it with other prominent automated ground truth creation methods. We also evaluate its downstream utility and its ability to generalize across different state-of-the-art LLMs. Finally, we evaluate the proposed HITL pipeline by comparing its impact on junior analysts, specifically in terms of how it influenced their precision, recall, and labeling speed.
@@ 5.1.1 Comparison with Existing Automated IoC Labeling Systems
We use PRISM, our dataset of IoCs that are manually labeled by multiple threat analysts (see Section 4), to evaluate the most prominent existing automated ground truth (GT) creation methods discussed in Section 2.2, and compare them with LANCE. Specifically, we assess four strategies based on their extraction and labeling performance: (1) IoC Searcher with whitelist-based filtering [1], (2) AlienVault [7], and (3,4) VirusTotal-based labeling with two thresholds [24, 35, 31, 30, 29]. None of the mentioned GT creation methods requires training.
Evaluation Setup: For the RegEx + Whitelist method, we implemented the GT creation pipeline from the GoodFATR paper [1], which supplements rule-based extraction by IoC Searcher with whitelist and frequency-based filtering to reduce false positives. The AlienVault method involved retrieving all IoCs (including inactive ones) linked to each report from the AlienVault platform [7]. All indicators not in the platform were considered unlabeled and treated like nonIoCs. For the VirusTotal-based methods, we followed prior work [30, 29, 24, 31] by extracting indicators with IoC Searcher and labeling them based on VirusTotal lookups. We used two thresholds: Threshold 1 labels an indicator as IoC if at least one vendor flags it as malicious, and as nonIoC if all vendors deem it benign. Threshold 5 raises the bar, requiring five or more malicious detections to label an indicator as IoC, while indicators with fewer than five malicious detections are labeled nonIoC. Indicators not included in the database are left unlabeled and are considered as nonIoC
Coverage Analysis: We assume a total of 1,789 candidate indicators, extracted using IoC Searcher, a state-of-the-art rule-based tool [1, 30, 24, 31]. Figure 6 shows the ratio of unlabeled indicators across methods. The RegEx + Whitelist method labels all extracted indicators but still suffers from poor precision due to insufficient context awareness. The AlienVault method yields the most unlabeled indicators (37%). When evaluating the ground truth provided by AlienVault, we also observe that the sum of Labeled and Unlabeled indicators exceeds the total of 1,789. This can be explained by the analysis in Section 2.3. The VirusTotal-based methods leave over 280 indicators unlabeled, suggesting that 16% of the indicators are absent from the VirusTotal database. In contrast, LANCE labeled over 99% of all extracted indicators, with the few unlabeled cases attributed to malformed LLM outputs.
@@
Performance Analysis: Figure 6 summarizes overall precision, recall, and F1 score for each method. We further break down performance across the four IoC types (IP, hash, URL, domain) in Appendix D (Table VII).
The RegEx + Whitelist method achieves almost perfect recall but low precision, particularly for domains and URLs, making it unsuitable for high-fidelity ground truth creation due to excessive false positives. AlienVault shows the lowest recall for URLs (25%) and significant gaps in domain coverage (73% recall), likely due to incomplete extraction. Its precision on hashes is also low (69%) because many labeled hashes do not appear in the report text.
VirusTotal with threshold 1 performs reasonably well, achieving an average F1 score of 86%, with consistent scores across indicator types. Threshold 5 has increased precision; however, it suffers from low recall due to strict filtering criteria, reducing its reliability for comprehensive dataset creation.
LANCE outperforms all other methods, consistently achieving over 90% F1 score across all types and 97.6% overall. Its lowest score, 86.8% recall on domains, is due to intentional disagreements between LANCE’s labels and the senior analyst in cases involving compromised legitimate websites. Notably, LANCE’s justifications showed awareness of this nuance, reinforcing the importance of contextual labeling aligned with downstream use cases. Appendix C reports some concrete examples of such scenarios. These discrepancies between LANCE and the senior analyst’s labels highlight an important challenge: the definition of what constitutes an IoC can, in some cases, be ambiguous as it may depend on factors such as the time of observation, the current threat landscape, and the specific downstream task where the IoC labels will be used (e.g., detection, attribution, or sharing)[30, 53, 54]. This further justifies the need for manual verification with a human-in-the-loop approach.
@@ 5.1.2 LANCE Comparison With Naive LLM Prompts
These results show that LANCE is critical in the pipeline. Rather than relying on naive prompting, LANCE strategically orchestrates LLM capabilities via prompt design and context-aware segmentation and achieves high precision on complex threat data. Without this pipeline, the LLM underperforms by 30% in F1 score (as shown in Figure 7).
Takeaway: Without structured guidance, LLMs like ChatGPT 4o underperform in IoC extraction and labeling. LANCE significantly boosts both precision and recall, underscoring the importance of a tailored pipeline.
@@ 5.1.3 Downstream Utility Evaluation
To evaluate the usefulness of the IoC labels automatically generated by LANCE, we use it to train and then measure the effect it had on the performance of IoCMiner [28], a Machine Learning-based IoC classification tool. IoCMiner was developed for labeling from tweets. Given that our dataset consists of threat reports, we adapt our data by extracting English sentences that include indicators. We only evaluate its performance on IPs, URLs, and hashes since it was designed to only extract these types.
Out of the 500 reports from ORKL mentioned in Section 4.2 we use 450 for training and the other 50, which are included in PRISM, for testing. To extract English sentences from the reports, we first split the text of the report on periods and then use IoC Searcher [1] to see if an indicator of the targeted type is included in each sentence. For the 450 reports of the training set, we generate labels for the indicator in each sentence using the two highest performing ground truth generation methods (VirusTotal with threshold=1 and threshold=5) as indicated in Figure 6, and LANCE, automatically generating three different training sets (VT1, VT5, and LANCE). We train IoCMiner [28] once for each different training set using the available code and without altering the hyperparameters, and evaluate the three separately trained versions of the model on PRISM. The result of the evaluation can be seen in Figure 8.
We observe that the model trained on the dataset labeled by LANCE achieves a superior F1 score in PRISM than the models trained on datasets labeled by the other two methods. This is consistent with previous evaluations of LANCE, as in Section 5.1.1 we see that LANCE has superior labeling capabilities than Virsus Total with thresholds 1 and 5 and better ground truth enables better-performing models [38, 39]. More specifically, we see that the LANCE-trained model outperforms the VT1-trained model by over 6% and the VT5-trained model by over 8% in terms of F1 score.
@@ 5.1.4 Generalization Across LLMs
We tested LANCE on four additional prominent LLMs: LLama 3.3 70b [56], Gemma 3 27b[57], Gemini 2.0 Flash [58], and Nvidia Llama 3.1 Nemotron 70b [59]. We test those LLMs on PRISM ground truth with the same prompts that we developed in section 3.2. For our experiments on GPT and Gemini, we used the provided APIs, while we used the open-source pre-trained models LLama, Nemotron, and Gemma that are available, with quantization, on 2 A40 GPUs with 46068 MB of memory each. To ensure a fair evaluation, we only tested the LLMs on the 25 reports of BAP (where the GPT-based LANCE labels were not available to analysts during ground truth creation), eliminating potential bias.
We evaluated the four LLMs on the BAP-generated part of the PRISM ground truth dataset. The results, summarized in Figure 9, show that Gemma and Gemini perform comparably to GPT, achieving total F1 scores of 0.98 and 0.92, respectively. Llama and Nvidia Nemotron demonstrated moderate performance, which could likely be improved with targeted prompt engineering. A more detailed analysis of this comparison is presented in Appendix B.
In terms of coverage, Gemini and Nvidia Nemotron failed to label only 19 and 22 indicators, representing 2.1% and 2.4% of the dataset, highlighting their strong labeling capability relative to other leading approaches (see Section 5.1.1). Llama and Gemma missed 84 (9.2%) and 102 (11.1%) indicators, respectively, indicating lower, though still substantial, labeling coverage.
In terms of processing time per report, GPT and Gemini required approximately 2–3 minutes, while Llama, Nemotron, and Gemma took around 20 minutes due to more limited computational resources. Precise timing comparisons are challenging, as external factors such as network latency and server response times can significantly influence the measurements and are not indicative of the models’ inherent performance.
@@ 5.2.1 LANCE’s Effect on Junior Analysts’ Labels
To evaluate the LANCE-generated and the junior analysts-generated labels, we compare the Precision, Recall, and F1 scores they achieved on the 25 reports from the Baseline Annotation Pass (BAP) part of the ground truth generation methodology. We evaluate the results based on the final ground truth labels of those 25 reports. In Figure 5, we see that LANCE performed on par with the junior analysts. In the more difficult indicator types, namely domains and URLs we see an increase of 1.9% and 3.7% respectively in F1 score. We also observed an increase of 1.4% in total Precision, with the same metric increasing by 5.3% in domains. This means that LANCE produced fewer false positive domains than the junior analysts. Finally, we can see that LANCE achieved 6.9% higher recall in URLs, producing fewer false negative URLs than the junior analysts.
@@
In Figure 5, we also see the performance the junior analysts and LANCE achieved on the reports from the Guided Annotation Pass (GAP) part of the ground truth creation. In this part, the junior analysts had access to the LANCE-generated labels and justifications during the labeling of the indicators. With the assistance and the additional information provided by LANCE, the junior analysts performed better in total F1 score than LANCE and, by extension, better than without LANCE assistance.

### v2

@@ 3.3 Custom User Interface
The user interface provides the analyst with the LANCE-generated label and justification. The analyst has the option to either accept or override the generated label. The interface also enables analysts to view justifications by hovering over indicators as well as adding additional comments for future reference. After the analyst finalizes all labels, they can export them in a structured format along with all justifications and additional comments optionally provided by the analyst.
@@ 4 PRISM Dataset Creation
We now describe how we use the HITL system described earlier (Figure 1) to create our PRISM dataset. The labeling pipeline we use is depicted in Figure 3. We start by describing our ground truth generation methodology with human vetting, and then discuss our concrete application of this methodology to the PRISM dataset. We describe PRISM first here, and later (in Section 5), use the manually verified IoC labels obtained during this dataset creation process to evaluate LANCE’s accuracy by measuring its level of agreement with expert threat analysts.
@@ 4.1 Methodology Overview
To further increase the quality of the PRISM dataset, as well as to evaluate the effects LANCE had on the junior analysts, we also employed a senior analyst. The senior analyst annotated only the reports that had indicators in which the junior analysts disagreed. This way, we minimized the time the senior analyst needed to spend on labeling indicators and used their expertise only on the difficult-to-label indicators.
@@ 4.1.1 Baseline Annotation Pass (BAP)
The reports that included at least one disputed indicator were tagged as disputed reports and were given to a senior analyst to determine the correct label of the indicators. These reports were given to the senior analyst through the same UI and with the same instructions given to the junior analysts.
The senior analyst was tasked to label all indicators in the disputed reports, not only the disputed indicators. This way, we could see if other indicators that were labeled falsely by both the junior analysts and the LLM.
@@ 4.1.2 Guided Annotation Pass (GAP)
For the second part of the ground truth creation setup, we used the second set of reports. The setup for GAP was similar to that of BAP, with the exception that the analysts did have access to the labels that were assigned to the indicators by LANCE, as well as the justifications that LANCE provided for this choice. This allowed the junior analysts to be “advised” by LANCE while labeling each indicator.
@@ 4.2 PRISM Dataset Details
To ensure relevance and quality, we filtered for reports that also appeared in AlienVault [12, 7], a widely used threat intelligence platform. Specifically, we selected reports published by AlienVault’s official user account, which aggregates high-quality reports from vendors such as Kaspersky, Palo Alto Networks, and Microsoft. From this filtered corpus, we identified 50 reports that (i) appeared on both ORKL and AlienVault, (ii) were fewer than 30 pages in length, and (iii) contained more than five unique indicators. Although this threshold provides only an upper bound on the number of true IoCs, it serves as an effective proxy for report richness. The 30-page limit was a practical choice to reduce the workload of human annotation during dataset creation.
@@
TABLE II: Analysis of non-unique indicator instances in reports for BAP and GAP, and in total
Disagreements between annotators, or between annotators and the LANCE-generated labels, were flagged as disputed. We identified 15 reports from BAP that had at least one disputed indicator and nine from GAP. These reports were escalated to the senior analyst, who re-labeled all indicators in those reports to ensure consistency and completeness. Figure 4 shows the agreement ratios between annotators and with LANCE.
@@
Figure 5: Comparison of LANCE and analyst performance across precision, recall, and F1 score during both annotation phases. The top row shows results from BAP, while the bottom row shows GAP.
@@ 5 IoC Labeling Evaluation
In this section, we evaluate LANCE by comparing it with other prominent automated ground truth creation methods. We also evaluate its downstream utility and its ability to generalize across different state-of-the-art LLMs. Additionally, we evaluate the proposed HITL pipeline by comparing its impact on junior analysts, specifically in terms of how it influenced their precision, recall, and labeling speed. Finally, we discuss some common cases of indicators that were misclassified by LANCE.
@@ 5.1.1 Comparison with Existing Automated IoC Labeling Systems
We use PRISM, our dataset of IoCs that are manually labeled by multiple threat analysts (see Section 4), to evaluate the most prominent existing automated ground truth (GT) creation methods discussed in Section 2.2, and compare them with LANCE. Specifically, we assess four strategies based on their extraction and labeling performance: (1) IoC Searcher with whitelist-based filtering [1], (2) AlienVault [7], and (3,4) VirusTotal-based labeling with two thresholds [24, 35, 31, 30, 29]. None of the mentioned GT creation methods require training.
Evaluation Setup: For the RegEx + Whitelist method, we implemented the GT creation pipeline from the GoodFATR paper [1], which supplements rule-based extraction by IoC Searcher with whitelist and frequency-based filtering to reduce false positives. The AlienVault method involved retrieving all IoCs (including inactive ones) linked to each report from the AlienVault platform [7]. All indicators not in the platform were considered unlabeled and treated like nonIoCs. For the VirusTotal-based methods, we followed prior work [30, 29, 24, 31] by extracting indicators with IoC Searcher and labeling them based on VirusTotal lookups. We used two thresholds: Threshold 1 labels an indicator as IoC if at least one vendor flags it as malicious, and as nonIoC if all vendors deem it benign. Threshold 5 raises the bar, requiring five or more malicious detections to label an indicator as IoC, while indicators with fewer than five malicious detections are labeled nonIoC. Indicators not included in the database are left unlabeled and are considered as nonIoC.
Coverage Analysis: We assume a total of 1,789 candidate indicators, extracted using IoC Searcher, a state-of-the-art rule-based tool [1, 30, 24, 31]. Figure 6 shows the ratio of unlabeled indicators across methods. The RegEx + Whitelist method labels all extracted indicators but still suffers from poor precision due to insufficient context awareness. The AlienVault method yields the most unlabeled indicators ($37\%$). When evaluating the ground truth provided by AlienVault, we also observe that the sum of Labeled and Unlabeled indicators exceeds the total of 1,789. This can be explained by the analysis in Section 2.3. The VirusTotal-based methods leave over 280 indicators unlabeled, suggesting that $16\%$ of the indicators are absent from the VirusTotal database. In contrast, LANCE labeled over $99\%$ of all extracted indicators, with the few unlabeled cases attributed to malformed LLM outputs.
@@
Performance Analysis: Figure 6 summarizes overall precision, recall, and F1 score for each method. We further break down performance across the four IoC types (IP, hash, URL, domain) in Appendix D (Table VIII).
The RegEx + Whitelist method achieves almost perfect recall but low precision, particularly for domains and URLs, making it unsuitable for high-fidelity ground truth creation due to excessive false positives. AlienVault shows the lowest recall for URLs ($25\%$) and significant gaps in domain coverage ($73\%$ recall), likely due to incomplete extraction. Its precision on hashes is also low ($69\%$) because many labeled hashes do not appear in the report text. VirusTotal with threshold 1 performs reasonably well, achieving an average F1 score of $86\%$, with consistent scores across indicator types. Threshold 5 has increased precision; however, it suffers from low recall due to strict filtering criteria, reducing its reliability for comprehensive dataset creation.
LANCE outperforms all other methods, consistently achieving over $90\%$ F1 score across all types and $97.6\%$ overall. Its lowest score, $86.8\%$ recall on domains, is due to intentional disagreements between LANCE’s labels and the senior analyst in cases involving compromised legitimate websites. Notably, LANCE’s justifications showed awareness of this nuance, reinforcing the importance of contextual labeling aligned with downstream use cases. We further analyze mislabeled examples in Section 5.3. These discrepancies between LANCE and the senior analyst’s labels highlight an important challenge: the definition of what constitutes an IoC can, in some cases, be ambiguous as it may depend on factors such as the time of observation, the current threat landscape, and the specific downstream task where the IoC labels will be used (e.g., detection, attribution, or sharing)[30, 56, 57]. This further justifies the need for manual verification with a human-in-the-loop approach.
@@ 5.1.2 LANCE Comparison With Naive LLM Prompts
These results show that LANCE is critical in the pipeline. Rather than relying on naive prompting, LANCE strategically orchestrates LLM capabilities via prompt design and context-aware segmentation and achieves high precision on complex threat data. Without this pipeline, the LLM underperforms by $30\%$ in F1 score (as shown in Figure 7).
Takeaway: Without structured guidance, LLMs like ChatGPT-4o underperform in IoC extraction and labeling. LANCE significantly boosts both precision and recall, underscoring the importance of a tailored pipeline.
@@ 5.1.3 Downstream Utility Evaluation
To evaluate the downstream utility of a LANCE-generated dataset, we use LANCE-generated labels to train a downstream IoC-extraction tool. For this evaluation, we chose IoCMiner [28] as the downstream IoC-extraction tool, as it was the only tool in Table I that requires training and had usable code. IoCMine [28] is a Machine Learning-based IoC classification tool developed for labeling IPs, URLs, and hashes from tweets. The format of the training dataset provided by the authors of IoCMiner is based on one-hot encoded bag-of-words representations, thus making the reconstruction of the original sentences infeasible. For this reason, we generate our datasets for this evaluation, using 450 reports that are separate and do not overlap with the 50 reports in PRISM. We use these reports to generate three different datasets, one with LANCE-generated labels and two generated by VirusTotal-based labels (with thresholds 1 and 5). We then use the three different datasets to train three different instances of the IoCMiner model and evaluate each of them on PRISM.
Specifically, to extract English sentences from the reports, we first split the text of the report on periods and then use IoC Searcher [1] to see if an indicator of the targeted type is included in each sentence. For the 450 reports of the training set, we generate labels for the indicator in each sentence using the two highest performing ground truth generation methods (VirusTotal with threshold=1 and threshold=5) as indicated in Figure 6, and LANCE, automatically generating three different training sets (VT1, VT5, and LANCE). We train IoCMiner [28] once for each different training set using the available code and without altering the hyperparameters, and evaluate the three separately trained instances of the model on PRISM. The result of the evaluation can be seen in Figure 8.
We observe that the model trained on the dataset labeled by LANCE achieves a superior F1 score in PRISM than the models trained on datasets labeled by the other two methods. This is consistent with previous evaluations of LANCE, as in Section 5.1.1 we see that LANCE has superior labeling capabilities than Virsus Total with thresholds 1 and 5 and better ground truth enables better-performing models [41, 42]. More specifically, we see that the LANCE-trained model outperforms the VT1-trained model by over $6\%$ and the VT5-trained model by over $8\%$ in terms of F1 score.
@@ 5.1.4 Generalization Across LLMs
We tested LANCE on four additional prominent LLMs: LLama 3.3 70b [59], Gemma 3 27b[60], Gemini 2.0 Flash [61], and Nvidia Llama 3.1 Nemotron 70b [62]. We test those LLMs on PRISM ground truth with the same prompts that we developed in Section 3.2. For our experiments on GPT and Gemini, we used the provided APIs, while we used the open-source pre-trained models LLama, Nemotron, and Gemma that are available, with quantization, on 2 A40 GPUs with 46 GB of memory each. To ensure a fair evaluation, we only tested the LLMs on the 25 reports of BAP (where the GPT-based LANCE labels were not available to analysts during ground truth creation), eliminating potential bias.
We evaluated the four LLMs on the BAP-generated part of the PRISM ground truth dataset. The results, summarized in Figure 9, show that Gemma and Gemini perform comparably to GPT, achieving total F1 scores of $0.98$ and $0.92$, respectively. Llama and Nvidia Nemotron demonstrated moderate performance, which could likely be improved with targeted prompt engineering. A more detailed analysis of this comparison is presented in Appendix C.
In terms of coverage, Gemini and Nvidia Nemotron failed to label only 19 and 22 indicators, representing $2.1\%$ and $2.4\%$ of the dataset, highlighting their strong labeling capability relative to other leading approaches (see Section 5.1.1). Llama and Gemma missed 84 ($9.2\%$) and 102 ($11.1\%$) indicators, respectively, indicating lower, though still substantial, labeling coverage.
In terms of processing time per report, GPT and Gemini required approximately 2–3 minutes, while Llama, Nemotron, and Gemma took around 20 minutes due to more limited computational resources. Precise timing comparisons are challenging, as external factors such as network latency and server response times can significantly influence the measurements and are not indicative of the models’ inherent performance. The performance of Llama, Nemotron, and Gemma is expected to be substantially faster on modern cloud deployments, considering our GPUs are 5 years old. We emphasize that these latency figures are implementation artifacts, not architectural barriers, and are not indicative of fundamental scalability limitations.
@@ 5.2.1 LANCE’s Effect on Junior Analysts’ Labels
To evaluate the LANCE-generated and the junior analysts-generated labels, we compare the Precision, Recall, and F1 scores they achieved on the 25 reports from the Baseline Annotation Pass (BAP) part of the ground truth generation methodology. We evaluate the results based on the final ground truth labels of those 25 reports. In Figure 5, we see that LANCE performed on par with the junior analysts. In the more difficult indicator types, namely domains and URLs, we see an increase of $1.9\%$ and $3.7\%$ respectively in F1 score. We also observed an increase of $1.4\%$ in total Precision, with the same metric increasing by $5.3\%$ in domains. This means that LANCE produced fewer false positive domains than the junior analysts. Finally, we can see that LANCE achieved $6.9\%$ higher recall in URLs, producing fewer false negative URLs than the junior analysts.
@@
In Figure 5, we also see the performance that the junior analysts and LANCE achieved on the reports from the Guided Annotation Pass (GAP) part of the ground truth creation. In this part, the junior analysts had access to the LANCE-generated labels and justifications during the labeling of the indicators. With the assistance and the additional information provided by LANCE, the junior analysts performed better in total F1 score than LANCE, and, by extension, better than without LANCE assistance. The observed drop in domain precision in GAP is due to a higher concentration of ambiguous, legitimate domains in the GAP subset. A more detailed analysis of those indicators can be found in Section 5.3.